Experiment Volume of acetic acid(mL) Moles of acetic acid mass of baking soda(g) Moles baking soda Total Moles Mole fraction baking soda Circumference(cm) Volume of carbon dioxide(cm3)
1 24 0.419 9.02 47.5
2 21 0.366 13.01 55
3 18 0.314 17.54 60.5
4 15 0.262 22.035 66.0
5 12 0.209 26.519 63.5
6 9 0.157 31.03 56.5

7

6 0.105 35.005 52.0
8 None None N/A N/A N/A 1.00 0.00 0.00

I need to know how to calculate and fill in the blank spaces above.
Answer #1

Moles of baking soda (n) can be determined as follows.

n = mass/molar mass

Here, molar mass of baking soda (NaHCO3) = 23+1+12+(3×16) = 84 g/mol

For example, in experiment 1, n = 9.02 g/(84 g/mol) = 0.10738 mol

Total moles = moles of acetic acid + moles of baking soda = 0.419 + 0.10728 = 0.52638 mol

Molefraction of baking soda = moles of baking soda/total moles = 0.10738 mol/0.52638 mol = 0.204

Volume of carbon dioxide = 0.10738 mol×22400 cm^3/(47.5 cm)^3 = 2405.312 cm^3 (since 1 mole of any gas ay STP occupies 22.4 L = 22400 cm^3 volume).